OpenStudy (anonymous):

Please help. One card is selected from an ordinary deck of 52. What is the probability as a common fraction for a heart and a diamond

OpenStudy (anonymous):

Do you know how many of the cards in a 52 card deck are black and how many are red?

OpenStudy (anonymous):

let A=black cards are drawn B=red card is drawn P(A)=26/52=1/2 similarly P(B)=? \[P \left( A \cup B \right)=P \left( A \right)+P \left( B \right)\]

OpenStudy (anonymous):

number of hearts=13 number of diamonds=13 total no.of cards=52 calculate as done in (a)

OpenStudy (anonymous):

Since all of the cards are either red or black, its somewhat of a silly question...assuming your worded it correctly. You're ALWAYS going to draw a red or black card when you draw one. How do you express a probability of something that is certain to happen every time?

OpenStudy (anonymous):

Right....well the question was worded "one card is selected from an ordinary deck of 52. Find the probability as a common fraction in its simplest form for each event: a) a black card or a red card b) a heart and a diamond

OpenStudy (anonymous):

Likewise, b seems a silly question. A card obviously cannot be a heart AND a diamond. It's impossible. How do you express the probability of something that will never happen?

OpenStudy (anonymous):

Are you seeing that the answer to a) is 1 and b) is 0 ?

OpenStudy (anonymous):

Yes..I see. thank you :)